Consider the probability that more than 94 out of 153 computers will not crash in a day. Assume the probability that a given computer will not crash in a day is 56 % .
Answer #1

Mean = 153*0.56 = 85.68

Standard deviation = = 6.14

Hence,

P(More than 94)

= P(X > 94.5) [Continuity Correction]

= P(z > (94.5 - 85.68)/6.14)

= P(z > 1.44)

= 0.0754